The position vector of a particle P of mass 6kg at time t seconds is r where r=[(18t-4t^3)I+ t^2j]m determine to 2 decimal places the value of t for which the velocity and the acceleration vectors of P are at right angles

1
Expert's answer
2021-11-24T17:18:28-0500

"\overrightarrow{r}" = [(18t-4t³)i+ t²j] m

Velocity will be

"\overrightarrow{v}=\frac{d\overrightarrow{r}}{dt}= [ (18 - 12t²)i + 2tj]" m/s

Acceleration will be

"\overrightarrow{a}=\frac{d\overrightarrow{v}}{dt}= [ (- 24t)i + 2j]" m/s²

Since velocity vector and acceleration vector are perpendicular "\overrightarrow{v}.\overrightarrow{a}" = 0

=> "[ (18 - 12t\u00b2)i + 2tj].[ (- 24t)i + 2j]=0"

=> [ -24t(18-12t²)+4t ]= 0

=> 4[-6t(18-12t²)+t] = 0

=> 72t³-108t+t = 0

=> 72t³ = 107t

=> 72t² = 107 as t ≠ 0

=> t = "\sqrt{\frac{107}{72}}"

=> t = 1.22 second [ correct up to 2 decimal places]